OpenStudy (anonymous):

Max has twice as much money as Katy, who has $12 more than Greg. All three together have $124. How much does each person have? ??? help?

OpenStudy (anonymous):

2m>k +12> g=124 ???

OpenStudy (anonymous):

M = 2 K K = G + 12 M + K + G = 124 Solve the above three equations

OpenStudy (anonymous):

let max's amount of money be 'x',katy's amount of money be 'y' and grey's amount of money be 'z' 'Max has twice as much money as Katy' x=2y----------1 'Katy has $12 more than Greg' y=z+12----------2 'All three together have $124' x+y+z=124--------3 now sove

OpenStudy (anonymous):

From 2, K - G = 12 Adding above and 3rd equation, we get M + 2K = 136 and from 1, M - 2K = 0 add this and the above, 2M = 136 M = 136/2 = 68 Therefore K = M / 2 = 68 / 2 = 34 and G= K-12 = 34 -12 = 22 Therefore, M=68, K=34 and G =22

OpenStudy (anonymous):

Thanks to both that was so helpful!!!